A particle moves in a circle of diameter 1.0 cm under the action of a magnetic field of 0.40 T. An electric field of 200 V/m makes the path straight. Find the charge/mass ratio of the particle.

Grade:10

1 Answers

Jitender Pal
9 years ago
Sol. r = 0.5 cm = 0.5 × 10^–2 m B = 0.4 T, E = 200 V/m The path will straighten, if qE = quB ⇒ E = rqB * B/m [∴ r = mv/qB] ⇒ E = rqB^2/m ⇒ q/m E/B^2r = 200/0.4 *0.4 * 0.5 * 10^-2 = 2.5 * 10^5 c/kg